Douglas bought a store and agreed to pay 100000 at the end of every six months for seven years. What is the equivalent cash price of the store if the interest rate is 5% compounded quarterly?
1
Expert's answer
2021-09-15T03:33:55-0400

Douglas bought a store and agreed to pay 100000 at the end of every six months for seven years

so each year he pays twice 100000 twice.

Amount he can pay for 7 years "=2\times 100000 \times 7=1400000"


If he opt to follow the second payment method, equivalent cash price of the store if the interest rate is 5% compounded quarterly can be calculated as follows.

"A=P(1+r)^n"

Since he is paying 100000 after every 6 months, so in a year he pays P=100000+100000=200000

", n=4\times 7=28, r=5/4=1.25\%"

"A=200000(1+1.25\/100)^{28}=283198.4607"